# On estimates of trace-norm distance between quantum Gaussian states

### v1

@@ 2 The algebra of CCR
$\alpha=\,\mathrm{diag}\left[\begin{array}[]{cc}a_{j}&0\\ 0&a_{j}\end{array}\right]_{j==1,\dots,s},$
@@ A Gaussian state $\rho_{m,\alpha}$ is determined by its quantum characteristic function
$\mathrm{Tr}\left(\hat{\alpha}-\hat{\beta}\right)\left(\hat{\beta}^{-1}-\hat{\alpha}^{-1}\right)$ (31)
@@ 3.2 Gauge-invariant states
implying that r.h.s. of (37) is upperbounded by
$\left\|\alpha^{2}-\beta^{2}\right\|_{1}$ $=$ $\left\|\left(\alpha-\beta\right)\circ\left(\alpha+\beta\right)\right\|_{1}$
$\leq$ $\left(\left\|\alpha\right\|+\left\|\beta\right\|\right)\left\|\alpha-\beta\right\|_{1}$
$=$ $\left(a+b\right)\left\|\alpha-\beta\right\|_{1},$
@@
$\left\|\left[2\left(A^{-1}B-I_{2s}\right)\beta\Upsilon_{\beta}\right]\circ\Upsilon_{\beta}\right\|_{1}$
$\leq$ $2\beta\left\|\left[2\left(A^{-1}B-I_{2s}\right)\Upsilon_{\beta}\right]\circ\Upsilon_{\beta}\right\|_{1,\beta}$
$\leq$ $4\beta\left\|A^{-1}B-I_{2s}\right\|_{1,\beta}.$
@@
From (25) this is evaluated as

### v2

@@ 2 The algebra of CCR
$\alpha=\,S^{t}\mathrm{diag}\left[\begin{array}[]{cc}a_{j}&0\\ 0&a_{j}\end{array}\right]_{j==1,\dots,s}S,$
@@ A Gaussian state $\rho_{m,\alpha}$ is determined by its quantum characteristic function
$\mathrm{Tr}\left(\hat{\alpha}-\hat{\beta}\right)\left(\hat{\beta}^{-1}-\hat{\alpha}^{-1}\right)$
$=$ $4\mathrm{Tr}[\left(\alpha-\beta\right)+\left(\alpha\Upsilon_{\alpha}-\beta\Upsilon_{\beta}\right)]\Delta^{-1}[\left(\alpha-\beta\right)-\left(\alpha\Upsilon_{\alpha}-\beta\Upsilon_{\beta}\right)]\Delta^{-1}$
@@ 3.2 Gauge-invariant states
implying that the second term in r.h.s. of (37) is upperbounded by
$4\left\|\alpha^{2}-\beta^{2}\right\|_{1}$ $=$ $4\left\|\left(\alpha-\beta\right)\circ\left(\alpha+\beta\right)\right\|_{1}$
$\leq$ $4\left(\left\|\alpha\right\|+\left\|\beta\right\|\right)\left\|\alpha-\beta\right\|_{1}$
$=$ $4\left(a+b\right)\left\|\alpha-\beta\right\|_{1},$
@@
$\left\|\left[2\left(A^{-1}B-I_{2s}\right)\Upsilon_{\beta}\right]\circ\Upsilon_{\beta}\right\|_{1}$
$\leq$ $2b\left\|\left[2\left(A^{-1}B-I_{2s}\right)\Upsilon_{\beta}\right]\circ\Upsilon_{\beta}\right\|_{1,\beta}$
$\leq$ $4b\left\|A^{-1}B-I_{2s}\right\|_{1,\beta}.$
@@
From (25) and (41) this is evaluated as
@@ 4 Discussion
Acknowledgment. The author is grateful to M.E.Shirokov for comments helping to improve the presentation.